OpenStudy (anonymous):

simplify √200x^7 would it be 4√50x^6

OpenStudy (chaise):

Hmm. I don't know what the answer would be. Let me solve it. \[\sqrt{(200x )^{7}}=200x^{7/2}\] I don't even know how you would go about simplifying this some more - I think this is as far a you can go, or if I went about it the right way. Is it (200x)^7 or 200(x^7)?

OpenStudy (anonymous):

the first one right after "Let me solve it" question: how do you get a nice square root symbol like that with superscript exponents and everything ..do you copy and paste from something?

OpenStudy (anonymous):

sorry.. i got that wrong... \[10*\sqrt{2*x ^{7}}\] just click equation.. then you can insert this stuff or learn the syntax.. it's pretty easy.. ^^

OpenStudy (chaise):

\[\sqrt{200x^7}\] \[=\sqrt{4\times25\times2}x^7\] \[=10\sqrt{2x^7}\] \[= 10(2x^{7/2})\] Use your equation editor at the bottom of your comment box, a little hard to use at first.

OpenStudy (anonymous):

oh equation.. got ya

OpenStudy (anonymous):

yeah that looks totally right on your last post thanks chaise

OpenStudy (chaise):

Pleasure :)

OpenStudy (anonymous):

10\[\sqrt{200x^7} = \sqrt{100*2*x^6*x} = 10*\sqrt{2}*x^3\sqrt{x}\]